# R&D Efforts in Cherenkov Imaging Technologies for Particle Identification in Future Experiments

### v1

@@ Abstract
Future upgrades of the ALICE and LHCb experiments at the CERN LHC, the ePIC experiment at the Electron–Ion Collider (EIC) at BNL, USA, and the PANDA and CBM experiments at FAIR are among the main drivers of R&D efforts in Cherenkov imaging technologies. However, the development of these technologies is not limited to these projects alone. Upgrades of AMBER at the CERN SPS, EIC-China, Belle II, ALADIN at CERN and the CLD detector at the CERN FCC are also making significant contributions to the advancement of Cherenkov imaging technologies. These leading future experiments share common challenges and solutions, which are being addressed within the newly formed DRD4 collaboration for detector R&D on particle identification (PID) and photon detectors [31]. DRD4 plays a central role in coordinating collaboration, fostering innovation, and advancing Cherenkov imaging technology. In this article, we review selected future experiments and upgrades of existing detectors that make extensive use of PID based on Cherenkov imaging; we will look into their physics objectives, the challenges imposed on PID technologies, the corresponding risk-mitigation strategies, and the associated R&D efforts, with particular emphasis on sensor technologies and optimization of radiator materials.
@@ 2.1 ALICE-3 upgrade
The proposed upgrade of the ALICE detector at CERN aims to realize a completely new experimental apparatus, ALICE3, for LHC Run 5 and beyond. This upgrade seeks to study the properties of the quark–gluon plasma in relativistic heavy-ion collisions with a level of precision and accuracy exceeding current limits [10]. Extensive PID capabilities are central to the ALICE3 upgrade. For example, the design target of the ALICE3 PID system is to achieve better than 3 e/, /K and K/p separation for momenta up to 2 GeV/c, 10 GeV/c, and 16 GeV/c, respectively, while providing acceptance over eight units of pseudorapidity () [46]. The nature of collider experiments limits the available space, and the high LHC luminosity results in a harshly crowded environment.
@@
To meet these PID requirements, the ALICE3 baseline design foresees the use of ultra-fast time-of-flight (TOF) and proximity focusing RICH detectors in both the barrel region (), referred to as bRICH, and the endcap region (), referred to as fRICH. Silicon photomultipliers (SiPMs) have been chosen as the baseline photo-sensors for the barrel region due to the presence of a solenoidal magnetic field of approximately 2 T. However, the choice of photo-sensors in the endcap region is more critical and remains subject to ongoing investigation and further research. The radiation level in the endcap regions is approximately two orders of magnitude higher (3.9 MeV neq/cm2) than in the barrel region (6.1 MeV neq/cm2). It has been observed that SiPM performance suffers severe limitations as the dark count rate (DCR) increases to tens of MHz/mm2 when the radiation level reaches MeV neq/cm2 [12, 24]. Under such conditions, commercial atomic-layer-deposition (ALD) coated MCP-PMTs [44] have been reported to operate without noticeable degradation. In addition, Large-Area Picosecond Photodetectors (LAPPDs) and High-Rate Picosecond Photodetectors (HRPPDs) are under consideration. These efforts highlight the synergy between photo-sensor development for the ePIC experiment and the ALICE3 upgrade.
@@
The use of SiPMs operated as TOF detectors has also been considered within the ALICE3 upgrade, motivated by promising R&D results. A thin layer of high-refractive-index material (SiO2, MgF2, or high-refractive-index glass) placed directly in front of the SiPM array allows the production of a large number of Cherenkov photons. Owing to the 1/ dependence of the time resolution, where is the number of detected photons, a timing resolution of approximately 20 ps is expected when five or more photons are detected. Given the excellent intrinsic performance of SiPMs, such performance is achievable and has been demonstrated in recent beam test studies for the ALICE3 upgrade. Furthermore, the use of multiple aerogel layers with different refractive indices as a radiator is under consideration to minimize emission point uncertainty and increase the number of detected photons. It is also being considered to fill the proximity gap with a C5F10O/N2 gas mixture, enabling operation of the bRICH as a threshold counter to veto electrons up to momenta of approximately 4 GeV/c. This option would require increasing the proximity gap between the radiator and the sensor from its nominal value of 20 cm to 35 cm. Further details can be found in [9].
@@
A dedicated beam test was performed [52] to validate the design principles of the bRICH at the CERN PS T10 beam test facility. The beam test used a setup with a hydrophobic aerogel tile as the radiator, produced by the Aerogel Factory Co. Ltd. (Chiba, Japan) [4], with a refractive index of 1.03. This choice of aerogel radiator is of synergistic interest with the ePIC experiment at the EIC. The setup was designed using two different types of SiPM sensors produced by Hamamatsu. Sensors from the S13361 series with 64 pixels were used to detect Cherenkov photons generated in a 1 mm thick quartz window acting as a secondary Cherenkov radiator, while sensors from the S13552 series with 128 channels were used to detect Cherenkov photons produced in the aerogel. In fig. 1 (a) and fig. 1 (b) we can see the set-up for their beam test. In fig. 1 (c), we can also notice the impact of 5 ns time gating. The results are promising and are in excellent agreement with expectations: (a) a Cherenkov angle of 242 mrad for saturated particles; (b) a single-photon angular resolution of 3.8 mrad; and (c) scaling with the number of detected photons allows the required and targeted bRICH angular resolution of 1.5 mrad to be achieved, along with substantial suppression of background hits when an appropriate timing window is applied. It is expected to couple the final ASIC design to annealed SiPMs operated at C after exposure to the radiation levels anticipated in the bRICH region [11].
@@ 2.2 LHCb upgrade
The PID system of the LHCb [14] currently comprises two RICH detectors to cover an unprecedented phase-space. RICH1 [21] is placed upstream of the magnet, and another named RICH2 [6] is placed downstream of the spectrometer magnet. The radiator gas in RICH1 has a Cherenkov threshold of 10 GeV/c for kaons and charged hadron PID is possible down to 2 GeV/c in the veto mode. The radiator gas in RICH2 provides good /K separation in the high momentum particle range up to 100 GeV/c [23]. The aim of the future upgrade of the LHCb PID sub-systems is to provide similar good performance in the HL-LHC era [47]. This would require a substantial improvement in both the precision of the measurement of the Cherenkov angle and the rejection of the background. Both the improvements can be achieved with larger number of detected photons. The advanced algorithm used in the RICH reconstruction, is capable to predict the Time of Arrival (TOA) of the Cherenkov photons within 10 ps of accuracy. It has been demonstrated that a substantial background can be rejected with a 5 ns time gating. Currently the photon detector chain for the RICH system consists of Multi-anode photomultiplier tubes (MaPMTs) read out by the CLARO ASICs on front-end board [17]. Between 2030 and 2033, the RICH detectors are expected to replace the CLARO ASIC with a fast ASIC known as FastRICH [33],[39] with no changes in the photo-senors. A configurable time gate reduces the TDC range to a configurable width up to 6.25 ns (with 25 ps TDC bins) or up to 25 ns (100 ps bins); therefore, reducing the outlier hits and requiring fewer bits per hit timestamp. The requirements for the future RICH readout ASIC are several: reduction of the bandwidth of the outgoing data, power consumption and radiation hardness, in addition to timing performance. The Upgrade II foresees the ASIC to withstand a fluence of approximately 2x1013 neq/cm2 and 12 kGy at the RICH1 sensor column, including a factor two safety margin on the predicted fluence. Furthermore, the ASIC should be compatible to lpGBT [48] and VL+ [66] radiation-hard chipset used for the LHCb optical readout. One can see the rejection of the amount of background in fig. 2 (a); in fig. 2 (b), we see the ASIC coupled to the MaPMT. In fig. 2 (d) and fig. 2 (e), one can see that the ASIC can be coupled to various types of photo-sensors, namely classical MaPMTs, SiPMs, and MCP-PMT-based LAPPDs. This brings us to understand the second part of the LHCb upgrade plan, namely during 2036 and beyond. On one hand, there is available time to carry on R&D activities to identify the optimal photo-sensor; on the other hand with increasing challenges to keep the single photon resolution as good as it is now. Therefore, considerations are made to quest for photocathode with high photon detection efficiency shifted towards larger wavelength. This allows minimizing the chromatic error and ensuring good single photon angular resolution. In fig. 2 c, we can see a comparison between the photon detection efficiency of classical PMTs and SiPM sensors with respect to the chromatic aberration of C4F10. Beams tests are conducted with classical PMTs, and SiPM as well as LAPPD like photo-sensors coupled to FastRICH ASIC to identify the optimal choice of photo-sensors (fig. 2 (d) and fig. 2 (e)). SiPM coupled to a micro-lens array are envisioned to improve granularity; such a concept is not totally new for the LHCb collaboration, as LHCb SciFi trackers are opting for this option [58]. Furthermore, the flat mirror of LHCb RICH-1, which reflects the photons to the sensor surface, has been considered in order to enter the acceptance and increase the curvature of the spherical mirrors (see fig. 2 (f)). Such an optical system ensures improved emission point uncertainty as well as reduces occupancy and pixel error [26]. However, this brings the flat mirror within the acceptance, and hence R&D efforts in this direction are assured.
@@
The gases with CnFn+2 structures are known as Saturated Fluro Carbons (SFC). Their non conductivity and optical properties (large refractive index and limited chromaticity) are key elements for their use as Cherenkov radiators. Their non-conductivity, non-flammability and radiation-resistance also make these gases excellent coolants. These gases, also currently in use by the LHCb RICH detectors, have high impact in the environment due to their high Global Warming Potential (GWP). The GWP of C4F10 is roughly 8500, similarly the GWP of CF4 is roughly 7000. In future, stricter policy on the usage of such gases imposes severe challenges in the operation of the detectors. LHCb primarily foresees the replacement of CF4 with CO2. However, this brings larger chromatic aberration and lower photon yield for RICH2, and the issue with RICH1 remains unsolved. Therefore, identification and validation of an alternative radiator gas are of extreme importance for LHCb. Newer spur-oxygenated fluoro-ketones, for example from the 3M NOVEC® range, with CnF2nO structures, can offer similar performance but with very low, or zero GWP. Nevertheless, several aspects like flammability, toxicity, radiation hardness, non-conductivity, optical properties are of extreme importance and therefore extensive R&D is ongoing in this direction [34]. Such an effort bring synergy to the quest for a similar solution by the ePIC experiment and the dual RICH detector in the Future Circular Collider (FCC) [36].
@@ 2.3 TORCH
Low momentum PID plays a critical role in reducing the systematic uncertainty by suppressing background of key measurements related to CP violation. The PID can be enhanced, particularly for low momentum hadrons, by the incorporation of an innovative 30 m2 area TOF detector in the LHCb setup to provide charged separation up to 10 (20) GeV/c momentum and beyond. To obtain such performance, the time of propagation of Cherenkov light in a fused silica window is measured with sensors of high timing performance. The TORCH (Time of internally Reflected CHerenkov light) detector would require a 15 ps timing resolution on a track by track basis to achieve a separation at 10 GeV/c. On one hand, such a detector would serve the role of the aerogel radiator of RICH1 that was envisaged in the original LHCb design, and on the other hand, it would serve as a general purpose timing detector [27, 35]. It has been reported that TORCH can provide 25 to 50% improvement in the tagging efficiency of initial state flavor tag, crucial for CP-violation measurements.
The TORCH aims to use planes of 1 cm thick fused silica as a source of prompt Cherenkov photons. It facilitates a modular design. Cherenkov photons travel to the periphery of the fused silica plates by total internal reflection, where they are reflected by a cylindrical mirror surface of a quartz block. This surface focuses the photons onto a plane of MCP-PMTs where their positions and arrival times are measured. The expectation is that typically 30 photons will be detected; hence, the required ToF resolution dictates the timing of single photons to a precision of around 70 ps. The MCP-PMTs have to withstand a high radiation environment (5 Ccm-2); therefore, the granularity and occupancy per pixel are subject to optimization. In the baseline design the TORCH are supposed to be operated with MCP-PMTs. The current generation of MCP-PMTs are known to have limitations in lifetime (due to ion feedback) and rate capability (from strip current), although improvements have been made, for example, through the use of ALD techniques. Improving the lifetime and rate capability of vacuum photon detectors is one of the goals of the CERN DRD4 collaboration. As an alternative, SiPMs are considered for the central regions of TORCH, where the occupancies are highest. SiPMs, however, have the drawback of large DCR, particularly when exposed to neutron fluences of 1013 neq /cm2 or more. In order to mitigate the DCR, it may be necessary to cool the SiPMs. Such a solution may adapt a common approach like the LHCb scintillating fibre detectors, where SiPMs are considered as a technology of choice. Currently, the FastRICH has been considered as the baseline for readout.

### v2

@@ Abstract
Future upgrades of the ALICE [4] and LHCb [25] experiments at the CERN LHC, the ePIC experiment at the Electron–Ion Collider (EIC) [1] at BNL, USA, and the PANDA [42] and CBM [3] experiments at FAIR are among the main drivers of R&D efforts in Cherenkov imaging technologies. However, the development of these technologies is not limited to these projects alone. EIC-China [16], upgrade of the Belle II [2], ALADDIN [11] at CERN and the CLD detector at the CERN FCC [17] are also making significant contributions to the advancement of Cherenkov imaging technologies. These leading future experiments share common challenges and solutions, which are being addressed within the newly formed DRD4 collaboration for detector R&D on particle identification (PID) and photon detectors [32]. DRD4 plays a central role in coordinating collaboration, fostering innovation, and advancing Cherenkov imaging technology. In this article, we review selected future experiments and upgrades of existing detectors that make extensive use of PID based on Cherenkov imaging; we will look into their physics objectives, the challenges imposed on PID technologies, the corresponding risk-mitigation strategies, and the associated R&D efforts, with particular emphasis on sensor technologies and optimization of radiator materials.
@@ 2.1 ALICE-3 upgrade
The proposed upgrade of the ALICE detector at CERN aims to realize a completely new experimental apparatus, ALICE3, for LHC Run 5 and beyond. This upgrade seeks to study the properties of the quark–gluon plasma in relativistic heavy-ion collisions with a level of precision and accuracy exceeding current limits. Extensive PID capabilities are central to the ALICE3 upgrade. For example, the design target of the ALICE3 PID system is to achieve better than 3 e/, /K and K/p separation for momenta up to 2 GeV/c, 10 GeV/c, and 16 GeV/c, respectively, while providing acceptance over eight units of pseudorapidity (). The nature of collider experiments limits the available space, and the high LHC luminosity results in a harshly crowded environment [46].
@@
To meet these PID requirements, the ALICE3 baseline design foresees the use of ultra-fast time-of-flight (TOF) and proximity focusing RICH detectors in both the barrel region (), referred to as bRICH, and the endcap region (), referred to as fRICH. Silicon photomultipliers (SiPMs) have been chosen as the baseline photo-sensors for the barrel region due to the presence of a solenoidal magnetic field of approximately 2 T. However, the choice of photo-sensors in the endcap region is more critical and remains subject to ongoing investigation and further research. The radiation level in the endcap regions is approximately two orders of magnitude higher (3.9 MeV neq/cm2) than in the barrel region (6.1 MeV neq/cm2). It has been observed that SiPM performance suffers severe limitations as the dark count rate (DCR) increases to tens of MHz/mm2 when the radiation level reaches MeV neq/cm2 [13, 26]. Commercial atomic-layer-deposition (ALD) coated MCP-PMTs have been reported to operate in high radiation levels without noticeable degradation; however, their performance has been noted to degrade above about 20 Ccm-2 [44]. In addition, Large-Area Picosecond Photodetectors (LAPPDs) and High-Rate Picosecond Photodetectors (HRPPDs) are also under consideration. These efforts highlight the synergy between photo-sensor development for the ePIC experiment and the ALICE3 upgrade. Presently, the choice of optimal photo-sensors for the endcap RICH remains an open question due to the level of radiation. Further studies and R&D efforts would be essential to identify the photo-sensors for the endcap RICH.
@@
The use of SiPMs operated as TOF detectors has also been considered within the ALICE3 upgrade, motivated by promising R&D results [52]. A thin layer of high-refractive-index material (SiO2, MgF2, or high-refractive-index glass) placed directly in front of the SiPM array allows the production of a large number of Cherenkov photons. Owing to the 1/ dependence of the time resolution, where is the number of detected photons, a timing resolution of approximately 20 ps is expected when five or more photons are detected. Given the excellent intrinsic performance of SiPMs, such performance is achievable and has been demonstrated in recent beam test studies for the ALICE3 upgrade [52]. Furthermore, the use of multiple aerogel layers with different refractive indices as a radiator is under consideration to minimize emission point uncertainty and increase the number of detected photons. It is also being considered to fill the proximity gap with a C5F10O/N2 gas mixture, enabling operation of the bRICH as a threshold counter to veto electrons up to momenta of approximately 4 GeV/c. This option would require increasing the proximity gap between the radiator and the sensor from its nominal value of 20 cm to 35 cm. Further details can be found in [5].
@@
A dedicated beam test was performed to validate the design principles of the bRICH at the CERN PS T10 beam test facility. The beam test used a setup with a hydrophobic aerogel tile as the radiator, produced by the Aerogel Factory Co. Ltd. (Chiba, Japan) [7], with a refractive index of 1.03. This choice of aerogel radiator is of synergistic interest with the ePIC experiment at the EIC. The setup was designed using two different types of SiPM sensors produced by Hamamatsu. Sensors from the S13361 series with 64 pixels were used to detect Cherenkov photons generated in a 1 mm thick quartz window acting as a secondary Cherenkov radiator, while sensors from the S13552 series with 128 channels were used to detect Cherenkov photons produced in the aerogel. In Fig. 1 (a) and Fig. 1 (b) we can see the set-up for their beam test. In Fig. 1 (c), we can also notice the impact of 5 ns time gating. The results are promising and are in excellent agreement with expectations: (1) a Cherenkov angle of 242 mrad for saturated particles; (2) a single-photon angular resolution of 3.8 mrad; and (3) scaling with the number of detected photons allows the required and targeted bRICH angular resolution of 1.5 mrad to be achieved, along with substantial suppression of background hits when an appropriate timing window is applied. It is expected to couple the final ASIC design to annealed SiPMs operated at C after exposure to the radiation levels anticipated in the bRICH region [12].
@@ 2.2 LHCb upgrade
The PID system of the LHCb [15] experiment currently comprises two RICH detectors to cover an unprecedented phase-space. RICH1 [24] is placed upstream of the magnet, and another named RICH2 [9] is placed downstream of the spectrometer magnet. The radiator gas in RICH1 has a Cherenkov threshold of 10 GeV/c for kaons and charged hadron PID is possible down to 2 GeV/c in the veto mode. The radiator gas in RICH2 provides good /K separation in the high momentum particle range up to 100 GeV/c [25]. The aim of the future upgrade of the LHCb PID sub-systems is to provide similar good performance in the HL-LHC era [47]. This would require a substantial improvement in both the precision of the measurement of the Cherenkov angle and the rejection of the background. Both the improvements can be achieved with larger number of detected photons. The advanced RICH reconstruction algorithm shows that the difference between the measured photon Time-of-Arrival (TOA) and the prediction from full LHCb detector simulation implies an intrinsic RICH timing resolution below 10 ps. A 5 ns timing gate is sufficient to reject a significant fraction of the background. Currently, the photon detector chain for the RICH system consists of Multi-anode photomultiplier tubes (MaPMTs) read out by the CLARO ASICs on front-end board [18]. For running between 2030 and 2033, the RICH detectors are expected to replace the CLARO ASIC with a fast ASIC known as FastRICH [35],[40] with no changes in the photo-sensors. A configurable time gate reduces the TDC range to a configurable width up to 6.25 ns (with 25 ps TDC bins) or up to 25 ns (100 ps bins); therefore, reducing the outlier hits and requiring fewer bits per hit timestamp. The requirements for the future RICH readout ASIC are several: reduction of the bandwidth of the outgoing data, power consumption and radiation hardness, in addition to timing performance. The Upgrade II foresees the ASIC to withstand a fluence of approximately 2x1013 neq/cm2 and 12 kGy at the RICH1 sensor column, including a factor two safety margin on the predicted fluence. Furthermore, the ASIC should be compatible to lpGBT [48] and VL+ [67] radiation-hard chipset used for the LHCb optical readout. In Fig. 2 (a), we see the ASIC coupled to the classical MaPMT. The ASIC is versatile and it can be coupled to various other types of photo-sensors, namely SiPMs, and MCP-PMT-based LAPPDs. This brings us to understand the second part of the LHCb upgrade plan, namely during 2036 and beyond. On one hand, there is available time to carry on R&D activities to identify the optimal photo-sensor; on the other hand with the increasing challenges to keep the single photon resolution as good as it is now. Therefore, the quest is for a photocathode with high photon detection efficiency shifted towards larger wavelength. This allows minimizing the chromatic error and ensuring good single photon angular resolution. Beam tests are conducted with classical MaPMTs, and SiPMs as well as LAPPD like photo-sensors coupled to FastRICH ASICs to identify the optimal choice of photo-sensors (Fig. 2 (b) and Fig. 2 (c)). SiPM coupled to a micro-lens array are envisioned to improve granularity; such a concept is not totally new for the LHCb collaboration, as LHCb SciFi trackers are opting for this option [58]. Furthermore, the flat mirror of LHCb RICH-1, which reflects the photons to the sensor surface, has been considered to enter the acceptance and increase the curvature of the spherical mirrors. Such an optical system ensures improved emission point uncertainty as well as reduces occupancy and pixel error [28]. However, this brings the flat mirror within the acceptance, and hence R&D efforts towards reducing the material budget are assured.
@@
The gases with CnFn+2 structures are known as Saturated Fluro Carbons (SFC). Their non-conductivity and optical properties (large refractive index and limited chromaticity) are key elements for their use as Cherenkov radiators. Their non-conductivity, non-flammability, and radiation-resistance also make these gases excellent coolants. These gases, also currently in use by the LHCb RICH detectors, have high impact in the environment due to their high Global Warming Potential (GWP). The GWP of C4F10 is roughly 8500, similarly, the GWP of CF4 is roughly 7000. In future, stricter policy on the usage of such gases imposes severe challenges in the operation of the detectors. LHCb primarily foresees the replacement of CF4 with CO2. However, this brings larger chromatic aberration and lower photon yield for RICH2, and the issue with RICH1 remains unsolved. Therefore, identification and validation of an alternative radiator gas are of extreme importance for LHCb. Newer spur-oxygenated fluoro-ketones, for example from the 3M NOVEC® range, with CnF2nO structures, can offer similar performance but with very low, or zero GWP. Nevertheless, several aspects like flammability, toxicity, radiation hardness, non-conductivity, optical properties are of extreme importance and therefore extensive R&D is ongoing in this direction [36]. Such an effort bring synergy to the quest for a similar solution by the ePIC experiment and the dual RICH detector in the Future Circular Collider (FCC) [38].
@@ 2.3 TORCH
Low momentum PID plays a critical role in reducing the systematic uncertainty by suppressing the background of key measurements related to CP violation. The PID can be enhanced, particularly for low momentum hadrons, by incorporating an innovative 30 m2 area TOF detector in the LHCb setup to provide charged separation up to 10 (20) GeV/c momentum and beyond. To obtain such performance, the time of propagation of Cherenkov light in a fused silica window is measured with sensors of high timing performance. The TORCH (Time of internally Reflected CHerenkov light) detector would require a 15 ps timing resolution on a track by track basis to achieve a separation at 10 GeV/c. On one hand, such a detector would serve the role of the aerogel radiator of RICH1 that was envisaged in the original LHCb design; on the other hand, it would serve as a general purpose timing detector [29, 37]. It has been reported that TORCH can provide a 25 to 50% improvement in the tagging efficiency of the initial state flavor tag, crucial for CP-violation measurements.
The TORCH aims to use planes of 1 cm thick fused silica as a source of prompt Cherenkov photons. It facilitates a modular design. Cherenkov photons travel to the periphery of the fused silica plates by total internal reflection, where they are reflected by a cylindrical mirror surface of a quartz block. This surface focuses the photons onto a plane of MCP-PMTs where their positions and arrival times are measured. The expectation is that typically 30 photons will be detected; hence, the required ToF resolution dictates the timing of single photons to a precision of around 70 ps. The MCP-PMTs have to withstand a high radiation environment beyond the current state of the art of 30 Ccm-2; therefore, the granularity and occupancy per pixel are subject to optimization. In the baseline design the TORCH is operated with MCP-PMTs. The current generation of MCP-PMTs are known to have limitations in lifetime (due to ion feedback) and rate capability (from strip current), although improvements have been made, for example, through the use of ALD techniques. Improving the lifetime and rate capability of vacuum photon detectors is one of the goals of the CERN DRD4 collaboration. As an alternative, SiPMs are considered for the central regions of TORCH, where the occupancies are highest. SiPMs, however, have the drawback of large DCR, particularly when exposed to neutron fluences of 1013 neq /cm2 or more. In order to mitigate the DCR, it may be necessary to cool the SiPMs. Such a solution may adapt a common approach like the LHCb scintillating fibre detectors, where SiPMs are considered as a technology of choice. Currently, the FastRICH has been considered as the baseline for readout.
In 2018, a test beam was conducted, demonstrating that a 70 ps single photon time resolution is achievable [20]. In 2025, a full-scale prototype was recently installed at the CERN PS beamline [45] and several critical aspects, such as mechanical design and optical gluing, have been validated.